If the perimeter of the window is 12 ft, what is the area A of the window as a function of the width x of the window?
Answer by scott8148(6628)

length plus width is half the perimeter

L + x = 6 ___ L = 6 - x

area is length times width

A = (6 - x) * x = 6x - x^2